OpenStudy (anonymous):

if it takes one person 40 mins to pick 3 quarts of berries at that rate how many minutes will it take 4 people to pick 15 quarts???

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

person takes 3/40 quart per minute ,4 people take 12/40 quart per min ,therefore to pick 15 quart they need 50 minutes

OpenStudy (anonymous):

the rate it takes for one person is 3/40 quarts per min "at that rate implies proportionality" so 4 people will be 4 times faster ie.\[4(3/40)=3/10\] \[3/10=15/x\] cross multiply \[3x=15*10\] x=50

OpenStudy (anonymous):

Thank you!!!!